# Everybody Prune Now: Structured Pruning of LLMs with Only Forward Passes

### v3

@@ 2.3 Selecting sub-models for evaluation
An as yet unexplored design choice is selecting the $n$ candidate sub-models for evaluation. A naive approach–uniform sampling–is shown here to be suboptimal. Take $m_{i}$ as a module that is critical for good performance under evaluation with $U$. Since $n<N$, it means that some modules may never be ”turned on” in the list of $n$ chosen sub-models. If $m_{i}$ happens to be one of these masked-out modules, the resulting estimate of $\hat{\beta_{i}}=0$ would in turn result in $\bm{m}^{\mathrm{approx}}$ being a poor estimate for $\bm{m}^{\ast}$. Thus, a more informed selection is necessary for accurate and useful $\bm{\beta}$ estimates.
@@
Given a module $m_{i}$, we set the likelihood of it being present in any of the $n$ sampled sub-models proportional to a prior $\rho_{i}$ which reflects its usefulness. To define $\rho_{i}$, we can turn to metrics from the pruning literature. Specifically, we set $\rho_{i}$ to be a module-level analogue of any of the pruning metrics like Wanda (Sun et al., 2023) or activation magnitude. For example, in one-hidden-layer network of dimension $d$, if activation magnitude is the prior, we compute the activation vector over multiple samples. The probability of retaining the $i$th column of $W\in R^{d\times d}$ is: $\rho_{i}\propto\hat{\bm{a}}_{i}=\frac{1}{B}\sum_{b}\big{|}\sigma\big{(}\big{(}W^{T}[i,:]\big{)}x_{b}\big{)}\big{|}$ where $\sigma\text{ is the nonlinearity}$. Sampling sub-models based on $\bm{\rho}$ biases selection toward high-performing models. Since $\rho_{i}$ can be computed via forward passes through the unmodified model $\bm{M}_{\theta}$, this approach remains memory-efficient. Appendix H details the priors we explored.
@@
2.4 Iterated Pruning
Previous work on gradient-based pruning (Anwar et al., 2017; Frankle & Carbin, 2018) have shown that taking an iterated approach to pruning yields improved results over pruning directly to the target sparsity $p$. Similarly, we define an updated pruning fraction $p_{\mathrm{iter}}<p$ and perform $\mathrm{iter}=\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ steps where we explore $n_{\mathrm{iter}}=\lceil\frac{n}{\mathrm{iter}}\rceil$ sub-models at a time. At the beginning of each iteration, we re-estimate the priors $\bm{\rho}$ for the unpruned modules and use them to select the $n_{\mathrm{iter}}$ sub-models to evaluate.
We combine the methods from Sections 2.2, 2.3, and 2.4 to develop Bonsai, a gradient-free structural pruning algorithm (Figure 2). Algorithm 1 provides the full details.
3 Experimental Details and Main Results
Table 1: Reported memory consumption of different methods. The minimum amount of memory required to run a Llama-7B model at half precision (FP16) is 14Gb. Running a forward pass with a batch size of 1 using the default model sequence length of 4096 uses around 20Gb of memory.
Forward Only Gradient-Based
Llama-2-7B (Model Only) Forward with bsz=1 (Bonsai Min.) Bonsai (Faster) LoRA Prune (Zhang et al., 2023) Compresso (Guo et al., 2023) LLM-Pruner (Ma et al., 2023) ShearedLlaMA (Xia et al., 2023)
14GB 20GB A6000 (48GB) 1 A100 (80GB) 4 V100 (128GB) 2 A100 (160GB) 8 A100 (640GB)
To reduce variance in score estimates, we average over 32 data points. A practitioner with 20GB of memory would run 32 forward passes with batch size 1, but this would be too slow for our experiments. Instead, we use a 48GB A6000 GPU (although not required by Bonsai), enabling batch sizes of 4–8 for faster experimentation.
@@ 3.1 Bonsai is competitive with other forward pass-only, structured pruning methods
Figure 3: Perplexity at 50% sparsity of LLaMA-{1,2}-7B on Wikitext-2 and C4 datasets.
@@
We focus our first set of experiments on comparing structured pruning methods that can be run without gradient-based optimization. We prune the LLaMA-2 7B model (Touvron et al., 2023) to 50% sparsity and evaluate on the Wikitext-2 (Merity et al., 2016) validation dataset. Our module importance signal for pruning, $U$, is the language modeling performance on the training set. When measuring speedups, we consider end-to-end latency of running inference on model.sequence_length chunks of the Wikitext-2 validation set. See Table 10 (Appendix C.3) for details about the hyper-parameters used for these experiments.
@@
Figure 1 shows the results of some of our experiments.  Bonsai expectedly outperforms the structured variant of Wanda (Sun et al., 2023), producing models with much lower perplexity at all sparsity levels and at any fixed desired speedup over the parent model. The more competitive baseline is FLAP (An et al., 2024), which was published a little over a month before the first release of our work. Even here (as shown in Table 3),  Bonsai outperforms FLAP while maintaining comparable speedups.
A key advantage of  Bonsai is its flexibility—it can incorporate various pruning methods as informative priors (magnitude, Wanda, or fluctuation-based metrics like those used in FLAP–see Appendix H for more details on how we estimate priors). Once these priors are established, Bonsai’s regression step refines the importance estimates, potentially yielding better pruning decisions.
@@
Bonsai offers practitioners flexible runtime-quality tradeoffs. While the reported results use a configuration that runs in $\leq~{}4hours$ (which is considerably more than FLAP’s $\approx 1$hr’s runtime), practitioners can configure  Bonsai  to produced structurally pruned models in as little as 15 minutes by reducing the number of perturbations, sample size, or pruning iterations—at at some cost to performance. Conversely, allocating more compute time generally yields better pruning outcomes. This flexibility allow practitioners to tailor Bonsai to their specific constraints, unlike methods with fixed computational profiles.
3.2 Introducing Post-Pruning Adaptation (PPA)
After pruning the parent model, depending on the sparsity level $p$ achieved, it is possible to obtain a pruned model on which we can run full fine-tuning or a parameter-efficient fine-tuning method like LoRA (Hu et al., 2021) with the available hardware memory. In this case, we can fine-tune the pruned model on the downstream task in order to recover more performance relative to that of the parent model.
$\mathcal{L}_{\mathrm{distill}}=\sum_{i}D_{\mathrm{KL}}\big{(}\mathrm{logits}^{i}\left(\mathbf{M}_{|\mathbf{m}^{\mathrm{approx}}}\right)~{}\|~{}\mathrm{logits}^{i}\left({\mathbf{M}}\right)\big{)}$, where $\lambda$ is a scalar weighting that can cross validated. Note, distillation can be performed without significant memory overhead by apriori caching the logits from the parent model $\mathbf{M}$ instead of hosting the model in memory during fine-tuning. In the sections ahead, we will apply PPA after pruning the parent model if the child model is small enough to allow for this.
Table 2: Wikitext-2 perplexity at 50% sparsity of LLaMA-2 7B with end-to-end latency speedups.
[…]
LlaMA-2 7B $5.11$ 1×1$\times$1 ×
Phi-2 3B $8.69$ 1.24×1.24$\times$1.24 ×
Wanda 2:4 3B $10.52$ 1.14×1.14$\times$1.14 ×
  + PPA $8.34$ 0.75$\times$
Bonsai 3B $19.47$ $\boldsymbol{1.58}\times$
  + PPA $8.89$ $\boldsymbol{1.58}\times$
3.2.1 Bonsai is competitive with semi-structured pruning methods
We compare Bonsai to the semi-structured variant of Wanda (Sun et al., 2023). In general, structured pruning methods under-perform semi-structured pruning methods, but compensate for this in speedup.
Before fine-tuning, the Wanda 2:4 model is more accurate but slower (1.14$\times$ vs 1.58$\times$) than the model from Bonsai. Since the Bonsai child model is small enough, we can perform PPA on it, resulting in improved accuracy (8.89ppl) with unchanged speedup.
Finetuning the semi-sparse Wanda 2:4 model is unfortunately less straightforward. It would require similar memory resources as needed to finetune the parent model222though the child tensors are sparse, the resulting gradients and cached intermediate tensors can be dense and have the same dimensions as those of the parent model (say $M\times M$). Since Bonsai does structured pruning, the actual tensor sizes are shrunk (say $N\times N~{}|~{}N<M$) which reduces memory during backward passes. but we are in the setting where we do not have enough memory for this. We therefore have to use a parameter efficient fine-tuning method like LoRA (Hu et al., 2021) instead. While the performance gap can be bridged by LoRA finetuning (10.52 $\rightarrow$ 8.34), the adapted semi-structured model experiences a drastic slowdown (0.75$\times$), since the learned low-rank matrices cannot be merged with the original sparsified ones without reverting back to dense computation. Thus, LoRA fine-tuned Wanda 2:4 is twice slower (0.48$\sim$ 0.48$\times$ 0.48 ×) than the model from Bonsai and similarly accurate.
In a memory-constrained setting, practitioners could opt for a pre-existing model of the target size instead of pruning a larger model. We compare the Bonsai-pruned model to Phi-2 (Li et al., 2023), a strong representative pre-existing model of similar size. As can be seen in Table 2, Bonsai is able to generate a model that is as accurate (0.2 difference in ppl) yet significantly faster (1.58$\times$ vs. 1.24$\times$ speedup), thus making it a competitive option to consider even if a model already exists at the target size.
3.2.2 Bonsai is competitive with gradient based structured pruning
Table 3: LLaMA-1 (50% sparsity) after post-pruning adaptation. are results as reported by (Zhang et al., 2023).
LLM-Pruner (Ma et al., 2023) 16.41 60.28 47.06 53.43 45.96 29.18 47.18
LoRAPrune (Zhang et al., 2023) 11.60 61.88 47.86 55.01 45.13 31.62 48.30
Bonsai + PPA 10.92 67.22 43.09 61.64 54.92 26.28 50.63
We compare Bonsai to the following gradient-based structured pruning approaches: LLM-Pruner (Ma et al., 2023) and LoRA-Prune Zhang et al. (2023). We use the reported results from (Zhang et al., 2023) since none of these methods are actually runnable in our memory-constrained setting (Table 1). We choose to compare to these over Sheared LlaMA (Xia et al., 2023) since they have much lower memory requirements (Table 1). We prune the LLaMA-1 7B model (Touvron et al., 2023) to 50% sparsity since these approaches report their results for the LLaMA-1 model only. We compare these methods on Wikitext-2 and also on six tasks from the Eleuther LLM Evaluation Harness benchmark (Gao et al., 2023). The pruning signal used for the Wikitext-2 task is the same as the above experiments. For the Eleuther Harness tasks, we use language modeling performance on the C4 (Raffel et al., 2020) dataset as pruning signal. We also perform parameter-efficient finetuning on our pruned model using 30K 512-length sequences from this corpus. Bonsai and LoRAPrune use similar amounts of the C4 dataset for Table 3 (30K vs 20K samples, respectively) whilst LLM-Pruner is trained on instruction tuned data with nearly twice the amount of unique samples (50K). Find more details in Appendix C.4.
As evinced by Table 3, Bonsai outperforms gradient-based methods even though it exclusively uses forward passes in the pruning stage. We attribute the superior performance of Bonsai to the fact that its pruning decisions are informed by directly exploring the space of sub-models whilst the other approaches resort to inaccurate proxies of module relevance in order to reduce the memory overhead of a fully gradient-based optimization approach (though not by enough to be runnable in our setting).
3.2.3 Bonsai can produce compressed models with strong zero-shot abilities
Generation Multiple Choice (MC)
(5-shot) (25-shot) (5-shot) (10-shot) (0-shot) (5-shot)
Phi-2 (Li et al., 2023) 2.7B 54.81 61.09 74.35 75.11 44.47 58.11 62.63
Phi-1.5 (Li et al., 2023) 1.5B 12.43 52.9 72.22 63.79 40.89 43.89 54.74
Sheared LlaMA (Xia et al., 2023) 1.3B Not Reported 33.5 57.9 60.7 Not Reported 25.7 *
Bonsai (w PPA) 1.8B 6.37 47.44 68.35 65.09 42.20 40.53 52.72
  + Reasoning Tuning 1.8B 27.67 45.56 68.82 64.51 42.58 40.97 52.49
Considerable amounts of compute and data, beyond what is feasible for lay ML practitioners, are needed to train LLMs with strong zero-shot capabilities (AI, 2023; Gemini Team et al., 2023). In this section, we demonstrate that Bonsai can empower everyday practitioners to produce strong and compact models with competitive zero-shot abilities by simply pruning bigger models on their available hardware.
We use Bonsai to prune a $\approx$3B Phi-2 model to $\approx$1.8B (35% sparsity). Bonsai hyper-parameters in this experiment are in Appendix C.5. Since it is small, the 1.8B pruned model can be fully fine-tuned on 1 A6000 GPU over 100k sequences of 2,048 tokens from the C4 dataset. As can be seen from Table 4, our pruned model achieves strong zero-shot performance on the Hugging Face OpenLLM leaderboard (Gao et al., 2023) compared to Phi-1.5, a smaller version in the Phi series of models that was trained from scratch.
Figure 4: LLaMA-2 7B @ 50% sparsity (No PPA). Pertubation and regression components are both needed to make Bonsai effective. Experiment details in Appendix D.
Interestingly, one exception to the general trend of Bonsai’s strong performance is the GSM-8K dataset, which is a mathematical reasoning dataset that requires generation of long reasoning chains. In our experiments, Bonsai prunes with respect to language modeling likelihood, as opposed to reasoning accuracy, and we posit that this mismatch may have contributed to our model’s underperformance on GSM8K. We attempt to remedy the drop in reasoning ability by including the GSM8K training data during post-pruning finetuning (resulting in a total of 108K fine-tuning samples). This boosts our model’s performance on the GSM8K with almost no degradation of performance on the other tasks.

### v4

@@ 3.3 Selecting sub-models for evaluation
An important design choice is selecting the $n$ candidate sub-models for evaluation. A naive approach such as uniform sampling would be suboptimal. Take $m_{i}$ as a module that is critical for good performance under evaluation with $U$. Since $n<N$, it means that some modules may never be "turned on" in the list of $n$ chosen sub-models. If $m_{i}$ happens to be one of these masked-out modules, the resulting estimate of $\hat{\beta_{i}}=0$ would in turn result in $\bm{m}^{\mathrm{approx}}$ being a poor estimate for $\bm{m}^{\ast}$. Thus, a more informed selection is necessary for accurate and useful $\bm{\beta}$ estimates.
@@
Given a module $m_{i}$, we instead set the likelihood of it being present in any of the $n$ sampled sub-models proportional to a prior $\rho_{i}$ which reflects its usefulness. To define $\rho_{i}$, we can turn to metrics from the pruning literature. Specifically, we set $\rho_{i}$ to be a module-level analogue of any of the pruning metrics like Wanda (Sun et al., 2023) or activation magnitude. For example, in one-hidden-layer network of dimension $d$, if activation magnitude is the prior, we compute the activation vector over multiple samples. The probability of retaining the $i$th column of $W\in R^{d\times d}$ is: $\rho_{i}\propto\hat{\bm{a}}_{i}=\frac{1}{B}\sum_{b}\big|\sigma\big(\big(W^{T}[i,:]\big)x_{b}\big)\big|$ where $\sigma\text{ is the nonlinearity}$. Sampling sub-models based on $\bm{\rho}$ biases selection toward high-performing models. Since $\rho_{i}$ can be computed via forward passes through the unmodified model $\bm{M}_{\theta}$, this approach remains memory-efficient. Appendix E details the priors we explored.
To enhance efficiency, we prune only the bottom $2p$ fraction of modules per layer, ranked by prior $\rho$, while keeping the top 12p1-2p fraction fixed.1112p2p is arbitrary; practitioners can tune it for optimal performance. This reduces the search space for sub-model evaluation. For the bottom $2p$ fraction, whenever we generate a mask $\alpha_{\bm{m}_{k}}$ with sparsity $p$, we also generate its complement $\alpha^{c}_{\bm{m}_{k}}$, obtained by flipping the values in $\alpha_{\bm{m}_{k}}$ (excluding the fixed 12p1-2p fraction). Covert and Lee (2020) show that this technique helps lower variance in regression with binary inputs.
3.4 Iterated Pruning
Previous works on gradient-based pruning (Anwar et al., 2017; Frankle and Carbin, 2018) have shown that taking an iterated approach to pruning yields improved results over pruning directly to the target sparsity $p$. Similarly, we define an updated pruning fraction $p_{\mathrm{iter}}<p$ and perform $\mathrm{iter}=\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ steps where we explore $n_{\mathrm{iter}}=\lceil\frac{n}{\mathrm{iter}}\rceil$ sub-models at a time. At the beginning of each iteration, we re-estimate the priors $\bm{\rho}$ for the unpruned modules and use them to select the $n_{\mathrm{iter}}$ sub-models to evaluate.
We combine the methods from Sections 3.2, 3.3, and 3.4 to develop Bonsai, a gradient-free structural pruning algorithm (Figure 2). Algorithm 1 provides the full details. A note about Line $5$ in our algorithm: depending on the task, we find that sampling $\mathbf{\bar{m}}_{i}$ and its complement $\mathbf{\bar{m}}^{c}_{i}$ helps reduce the variance of our regression estimate and leads to better results.
4 Experimental Details and Main Results
Table 2: Reported memory consumption of different methods. The minimum amount of memory required to run a LlaMA-7B model at half precision (FP16) is 14GB. Running a forward pass with batch size of 1 using the default model sequence length of 4096 uses around 20GB of memory.
Forward Only Gradient-Based
Base Model Bonsai (Min.) Bonsai (Faster) LoRA Prune Compresso LLM-Pruner Sheared LlaMA
LlaMA-2-7B Forward + Bsz=1 (Zhang et al., 2024) (Guo et al., 2023) (Ma et al., 2023) (Xia et al., 2024)
14GB 20GB 48GB 80GB 128GB 160GB 640GB
FP16 only A6000 1×A6000 1×A100 4×V100 2×A100 8×A100
In this section, we empirically evaluate Bonsai across multiple structured pruning settings. We compare Bonsai + PPA against both semi-structured pruning (Section 4.1) and gradient-based pruning (Section 4.2). Next, we demonstrate that Bonsai can yield compact models with strong zero-shot abilities (Section 4.3). Finally, we compare Bonsai to other forward-pass-only pruning methods (Section 4.4) to establish competitiveness under memory constraints.
To reduce variance in score estimates, we average over 32 data points. A practitioner with 20GB of memory would run 32 forward passes with batch size 1, but this extends experiment runtime. Instead, we use a 48GB A6000 GPU (although not required by Bonsai), enabling batch sizes of 4–8 for faster experimentation.
Introducing Post-Pruning Adaptation (PPA).
A central strength of Bonsai is that, by reducing models to a size that fits on commodity inference hardware depending on the sparsity level $p$ achieved, it unlocks the ability to further fine-tune (full fine-tuning or a parameter-efficient fine-tuning method like LoRA (Hu et al., 2021)) the pruned model on the same hardware. PPA is not required for Bonsai to be effective; but it shows that Bonsai does more than prune, since it expands access to to the full model adaptation pipeline in settings where gradient-based pruning is infeasible from the start.
$\mathcal{L}_{\mathrm{distill}}=\sum_{i}D_{\mathrm{KL}}\big(\mathrm{logits}^{i}\left(\mathbf{M}_{|\mathbf{m}^{\mathrm{approx}}}\right)~\|~\mathrm{logits}^{i}\left({\mathbf{M}}\right)\big)$, where $\lambda$ is a scalar weighting that can be cross-validated. Note, distillation can be performed without significant memory overhead by a priori caching the logits from the parent model $\mathbf{M}$ instead of hosting the model in memory during fine-tuning. In the subsections below, we will apply PPA after pruning the parent model if the child model is small enough to allow for this.
Table 3: Pruning can match highly optimized small, pretrained models while delivering better speedups. Bonsai-pruned LLaMA-2 achieves accuracy comparable to Phi-2 (a carefully engineered 3B model) with superior inference efficiency. Semi-structured pruning loses speedup benefits due to incompatible sparsity patterns.
@@ 4 Experimental Details and Main Results
LlaMA-2 7B $5.11$ $1\times$
Phi-2 3B $8.69$ $1.24\times$
Wanda 2:4 3B $10.52$ $1.14\times$
[…]
Bonsai 3B $19.47$ $\bm{1.58}\times$
  + PPA $8.89$ $\bm{1.58}\times$
4.1 Bonsai is competitive with semi-structured pruning methods
We compare Bonsai to the semi-structured variant of Wanda (Sun et al., 2023). In general, structured pruning under-perform semi-structured pruning, but compensate for this in speedup.
[…]
4.2 Bonsai is competitive with gradient-based structured pruning
[…]
LLM-Pruner (Ma et al., 2023) 16.41 60.28 47.06 53.43 45.96 29.18 47.18
LoRAPrune (Zhang et al., 2024) 11.60 61.88 47.86 55.01 45.13 31.62 48.30
Bonsai + PPA 10.92 67.22 43.09 61.64 54.92 26.28 50.63
Next we compare Bonsai to the following gradient-based structured pruning approaches: LLM-Pruner (Ma et al., 2023) and LoRA-Prune Zhang et al. (2024). We use the reported results from Zhang et al. (2024) since none of these methods are runnable in our memory-constrained setting (Table 2). We choose to compare to these over Sheared LlaMA (Xia et al., 2024) since they have much lower memory requirements (Table 2). We prune the LlaMA-1 7B model (Touvron et al., 2023) to 50% sparsity since these approaches report their results for the LlaMA-1 model only. We compare these methods on Wikitext-2 and also on six tasks from the Eleuther LLM Evaluation Harness benchmark (Gao et al., 2023). The pruning signal used for the Wikitext-2 task is the same as the above experiments. For the Eleuther Harness tasks, we use language modeling performance on the C4 (Raffel et al., 2020) dataset as pruning signal. We also perform parameter-efficient fine-tuning on our pruned model using 30K 512-length sequences from this corpus. Bonsai and LoRAPrune use similar amounts of the C4 dataset for Table 4 (30K vs 20K samples, respectively) whilst LLM-Pruner is trained on instruction tuned data with nearly twice the amount of unique samples (50K). Find more details in Appendix A.3.
As seen in Table 4, Bonsai outperforms gradient-based methods even though it exclusively uses forward passes in the pruning stage. We attribute the superior performance of Bonsai to the fact that its pruning decisions are informed by directly exploring the space of sub-models whilst the other approaches resort to inaccurate proxies of module relevance in order to reduce the memory overhead of a fully gradient-based optimization approach (though not by enough to be runnable in our setting).
4.3 Bonsai can produce compressed models with strong zero-shot abilities
Generation Multiple Choice (MC)
(5-shot) (25-shot) (5-shot) (10-shot) (0-shot) (5-shot)
Phi-2 (Li et al., 2023) 2.7B 54.81 61.09 74.35 75.11 44.47 58.11 62.63
Phi-1.5 (Li et al., 2023) 1.5B 12.43 52.9 72.22 63.79 40.89 43.89 54.74
Sheared LlaMA (Xia et al., 2024) 1.3B Not Reported 33.5 57.9 60.7 Not Reported 25.7 *
Bonsai (w PPA) 1.8B 6.37 47.44 68.35 65.09 42.20 40.53 52.72
  + Reasoning Tuning 1.8B 27.67 45.56 68.82 64.51 42.58 40.97 52.49
Considerable amounts of compute and data, beyond what is feasible for many practitioners, are needed to train LLMs with strong zero-shot capabilities (AI, 2023; Gemini Team et al., 2023). In this section, we demonstrate that Bonsai can empower everyday practitioners to produce strong and compact models with competitive zero-shot abilities by simply pruning bigger models on their available hardware.
We use Bonsai to prune a $\approx$3B Phi-2 model to $\approx$1.8B (35% sparsity). Bonsai hyper-parameters in this experiment are in Appendix A.4. Since it is small, the 1.8B pruned model can be fully fine-tuned on 1 A6000 GPU over 100k sequences of 2,048 tokens from the C4 dataset. As can be seen from Table 5, our pruned model achieves strong zero-shot performance on the Hugging Face OpenLLM leaderboard (Gao et al., 2023) compared to Phi-1.5, a smaller version in the Phi series of models that was trained from scratch.
Interestingly, one exception to the general trend of Bonsai’s strong performance is GSM8K, a mathematical reasoning dataset that requires generation of long reasoning chains. In our experiments, GSM8K performance is initially lower, consistent with task-agnostic pruning behavior in prior work (Xia et al., 2024; Reda et al., 2025); pruning based on language modeling can deprioritize modules needed for specialized reasoning. We attempt to remedy the drop in reasoning ability by adding 8K GSM8K samples during PPA (resulting in a total of 108K fine-tuning samples). This boosts our model’s performance on the GSM8K with almost no degradation of performance on the other tasks.
@@ 4.4 Bonsai is competitive with other forward pass-only, structured pruning methods
Table 6: Perplexity at 50% sparsity of LlaMA-{1,2}-7B on Wikitext-2 and C4.
@@
We focus our last set of experiments on comparing structured pruning methods that can be run without gradient-based optimization. We prune the LlaMA-2 7B model (Touvron et al., 2023) to 50% sparsity and evaluate on the Wikitext-2 (Merity et al., 2016) validation dataset. Our module importance signal for pruning, $U$, is the language modeling performance on the training set. When measuring speedups, we consider end-to-end latency of running inference on model.sequence_length chunks of the Wikitext-2 validation set. See Table 11 (Appendix A.2) for details about the hyper-parameters used for these experiments.
@@
Figure 1 shows that at 50% sparsity, Bonsai achieves the lowest perplexity among methods accessible in memory-constrained settings ($\leq 48$GB).  Bonsai expectedly outperforms the structured variant of Wanda (Sun et al., 2023), producing models with much lower perplexity at any fixed desired speedup over the parent model. A more competitive baseline is FLAP (An et al., 2024), which can be seen as a special case of our more general framework; here (as shown in Table 6),  Bonsai outperforms FLAP while maintaining comparable speedups.
@@
A key advantage of  Bonsai is its flexibility—it can incorporate various pruning methods as informative priors. In practice, we explored three classes of priors: activation magnitude (measuring the average size of activations per module), Wanda (a weight-activation product metric shown to be effective in unstructured pruing), and fluctuation-based metrics (as in FLAP, which track statistical variation across activations. These priors provide a a quick, forward-pass-only signal of which modules are likely to be important, and Bonsai’s regression step then refines them into more accurate importance estimates. See Appendix E for full details on how we estimate priors. An important outcome of this is that Bonsai offers practitioners flexible runtime-quality tradeoffs. While the reported results use a configuration that runs in $\leq~4hours$ (which is considerably more than FLAP’s $\approx 1$ hour runtime), practitioners can configure  Bonsai  to produce structurally pruned models in as little as 15 minutes by reducing the number of perturbations, sample size, or pruning iterations, potentially at some cost to performance. Conversely, allocating more compute time generally yields better pruning outcomes. This flexibility allows practitioners to tailor Bonsai to their specific constraints, unlike methods with fixed computational profiles.